# ChartAssisstant: A Universal Chart Multimodal Language Model via Chart-to-Table Pre-training and Multitask Instruction Tuning

### 10 Jan 2024 v2

@@ Abstract
Charts play a vital role in data visualization, understanding data patterns, and informed decision-making. However, their unique combination of graphical elements (e.g., bars, lines) and textual components (e.g., labels, legends) poses challenges for general-purpose multimodal models. While vision-language models trained on chart data excel in comprehension, they struggle with generalization and require task-specific fine-tuning. To address these challenges, we propose ChartAssistant, a chart-based vision-language model for universal chart comprehension and reasoning. ChartAssistant leverages ChartSFT, a comprehensive dataset covering diverse chart-related tasks with basic and specialized chart types. It undergoes a two-stage training process, starting with pre-training on chart-to-table parsing to align chart and text, followed by multitask instruction-following fine-tuning. This approach enables ChartAssistant to achieve competitive performance across various chart tasks without task-specific fine-tuning. Experimental results demonstrate significant performance gains over the state-of-the-art UniChart method, outperforming OpenAI’s GPT-4V(ision) on real-world chart data. The code and data are available at https://github.com/OpenGVLab/ChartAst.
@@ 1 Introduction
Figure 1: A comparison between previous chart-based models and our proposed ChartAssistant. ChartAssistant first aligns the chart and the text by pre-training on the chart-to-table translation task. After performing multitask instruction tuning, it can solve various downstream tasks without task-specific fine-tuning procedures.
@@
However, chart comprehension is challenging due to the intricate visual marks (e.g. lines, bars and symbols), implicit numerical information, and complex spatial relationships between elements (e.g. axes and labels). Interpreting charts requires specialized knowledge, spatial reasoning, and numerical understanding. The advanced general-purpose multimodal models [49, 21, 48] such as GPT-4V(ision) [3], trained on natural images, struggle with chart-related tasks due to the specific complexities and relationships unique to charts. Although recent multimodal literate models [29, 19] have achieved impressive results in processing various document-level tasks, they still face difficulties in accurately answering chart-related questions.
@@
In pursuit of universal chart reasoning and comprehension, prior works propose pre-training vision-language models on chart-related tasks as shown in Fig.1(a). For example, both Matcha [25] and UniChart [31] undergo multitask instructional tuning and task-specific fine-tuning, exhibiting good performance on several downstream tasks. However, these models still have severe downsides. Firstly, prior models fail to explicitly align the chart and the associated structured text-form table, which is essential to interpret the relationships between elements in the chart. Secondly, the existing training data [32, 30] is deficient in image-text annotations aimed at improving the model’s comprehension of visual elements and mathematical reasoning, as well as annotated data from the specialized chart types such as box-plots. Due to the above factors, existing chart-based models have poor generalization and require task-specific fine-tuning to achieve promising results on various downstream tasks as illustrated in Fig.1(a).
@@
To address these challenges, we propose ChartAssistant, a new vision-language model for universal chart comprehension and reasoning. We implement ChartAssistant with two variants, i.e. ChartAst-D and ChartAst-S, ChartAst-D is built upon Donut [16], a powerful vision-language model for visual document understanding. It employs Swin Transformer [28] as the visual encoder to process high-resolution charts and Bart [20] decoder to produce the expected output conditioned on the encoded image and text. While ChartAst-S is built upon SPHINX [23], a large vision-language model for universal multimodal comprehension. SPHINX employs multiple visual encoders for mixed encoding and leverages the powerful language capabilities of LLM [49]. Additionally, it incorporates specialized processing techniques for high-resolution images, resulting in enhanced chart comprehension abilities. Compared to ChartAst-D, ChartAst-S is a more advanced version, offering increased robustness and usability. It excels in its ability to handle a wide range of tasks and exhibits strong performance in various scenarios. To improve generalization, ChartAssistant is trained on a large-scale chart-specific benchmark dubbed ChartSFT. The training process involves a two-stage pre-training pipeline which employs chart-to-table pre-training to align the chart and its structured text and then perform joint tuning on multiple chart-related tasks as shown in Fig.1(b).
@@
Before conducting multitask instruction tuning, as done in existing research [31, 25], we begin by pre-training ChartAssistant on the chart-to-table translation task as shown in Fig.1(b). This task involves parsing a chart and generating a Markdown table. It shares similarities with dense captioning for natural images, allowing the model to interpret the elements and relationships within the chart. Similar to the role of image captioning in training multimodal models [27, 38, 45], chart-to-table translation facilitates alignment between the chart and its structured text. Following pre-training, we proceed with multitask instruction tuning using ChartSFT. This two-stage training approach enables ChartAssistant to achieve strong performance across a range of chart-related tasks without requiring task-specific fine-tuning.
The contributions of this paper can be summarized as follows. 1) We present ChartAssistant, a vision-language model for chart comprehension and reasoning. ChartAssistant is versatile enough to solve various chart-related tasks across a wide range of chart types. 2) We build a chart-specific visual instruction-following benchmark dubbed ChartSFT. ChartSFT surpasses existing chart-based benchmarks with its larger instruction-following data corpus, a broader range of tasks and chart types, and more comprehensive data annotations. 3) Extensive experimental results on various downstream tasks demonstrate that ChartAssistant surpasses the previous SoTA method UniChart [31] by 210.0%, 479.0%, 15.4% performance gain on numerical QA, referring QA and ChartQA, respectively. Notably, ChartAssistant even outperforms advanced multimodal models such as OpenAI’s GPT-4V(ision) [3] and Google’s Bard [6] on chart data in real scenarios.
@@ 2.2 Chart-specific Vision-Language Model
Some methods modify vision-language models for chart-related tasks [7, 26] or develop plugin for LLM to understand the chart [44]. Matcha [25] extends Pix2Struct [19] by integrating mathematical reasoning and chart data extraction tasks, excelling at chart question answering and chart summarization. Unichart [31] undergoes multitask instruction tuning for more chart-related tasks, establishing itself as the most versatile and effective chart vision-language model currently available. However, these methods have limitations. They require task-specific fine-tuning for promising results, risking catastrophic forgetting [46, 5] and sacrificed task generalization. Furthermore, these models struggle with mathematical computations, limiting their effectiveness and range of applicable chart types.
Contrastingly, we propose ChartSFT, the most extensive dataset to date, supporting a wide variety of chart tasks and types. We develop ChartAssistant using ChartSFT with a two-stage training strategy, capable of handling diverse chart-related tasks without requiring task-specific fine-tuning.
@@ 3 ChartSFT
We construct a large-scale chart-specific benchmark called ChartSFT by collecting instruction-following data from various tasks. The composition of ChartSFT is shown in Table 1, as extensively described below. Our ChartSFT consists of $39388322$ pieces of chart-text annotated data, $4.75$ and $5.62$ times larger than MatCha [25] and UniChart [31], respectively, as illustrated in Fig.3. ChartSFT contains chart types including both base and specialized types. Specifically, our ChartSFT encompasses nine types of charts by collecting data from various sources as shown in table 8. First, most charts with base types including bar, line, dot-line, and pie are collected from several existing datasets [30, 32, 14, 37, 22, 39, 15]. Second, we also generate some charts with base types from arxiv tables [1] and data augmentation techniques (e.g. various APIs and figure parameters). In particular, we use ChatGPT to suggest the proper chart type given each table data from arxiv. Third, we synthesize table data which is appropriate for depicting charts with specialized types.
@@ 3.2 Charts with Specialized Types
Previous Chart-based models perform poorly on charts with specialized types including radar, area, histogram, bubble, and box-plot. To improve generalization, we also train our ChartAssistant on charts with these specialized types. We collect chart data on specialized types through the following pipeline: i) pairs of chart and table generation, and ii) multitask dataset generation.
Chart-to-Table data generation. Considering the significant challenge of collecting real-world data on a large scale, we generate data synthetically. The (chart, table) pairs are generated in the following steps:
i) Table generation: We randomly select one distribution from over twenty predefined probability density distributions, and then generate tabular data by utilizing this distribution for randomization. For radar, bubble and area charts, We directly utilize randomly generated numbers as the table data. For histogram and box plot, we generate an array of extensive numbers using this distribution and calculate the statistical metrics of this array to serve as the table data (e.g., frequencies corresponding to histograms, upper whiskers corresponding to box plots). And then we use the generated data to prompt ChatGPT for creating titles, legends, and labels that align with the numerical characteristics.
ii) Chart generation: We utilize multiple Python plotting libraries to generate a variety of charts, including styles from matplotlib, plotly, pyecharts, ggplot, seaborn, altair, and more. To ensure diversity in the charts, we randomly select the following parameters: line style, font style, colors, markers, the position of legends and so on. Besides our own synthetic data, we also use the table from PlotQA [32], ChartQA [30], ChartSumm [37] and Chart-To-Text [15] to plot the charts for area and radar charts.
Multitask dataset generation. For the chart summarization and open-ended QA tasks, we instruct ChatGPT to build datasets by supplying both the table and the corresponding types of charts. For numerical QA and referring QA tasks, we adhere to the approach of the chart with base types by crafting a series of mathematical question templates tailored to the distinct characteristics of various chart types. Subsequently, we manually generate answers with COT annotations.
@@ 4.1 Architecture
The key to completing the tasks related to charts lies in accurately understanding the content of the charts. As shown in Fig. 4, we implement ChartAssistant with two variants, i.e. ChartAst-D and ChartAst-S, which have 260M and 13B parameters in total, respectively. Both ChartAst-D and ChartAst-S perform well in many chart-related tasks but ChartAst-D has a smaller size and ChartAst-S enjoys better generalization.
@@
ChartAst-S is a large vision-language model for chart understanding built upon Sphinx [23]. For high-resolution images, it preserves the original information through sampling and partitioning methods, ensuring greater fidelity to the image content. Moreover, Sphinx leverages the abundant prior knowledge of LLM [40] to handle various tasks such as visual question answering and image summarization. Specifically, for an input image $X_{V}$. ChartAst-S incorporates multiple visual encoders to extract more informative visual features $Z_{V}$, such as DINOv2 [33], CLIP [35], and ConvNeXt [43]. Unlike ChartAst-D where visual tokens are involved in a language decoder with a cross-attention module, ChartAst-S directly appends visual tokens to the text tokens $X_{q}$. The merged tokens are then fed into the LLM to generate the response. Thanks to the intricate design of the visual encoder and the powerful reasoning ability of LLM, ChartAst-D performs well in various real-world chart-related applications.
@@ 4.2 Training
In our ChartSFT, we have a corresponding instruction $X_{q}$ and response $Y_{q}$ for each image $X_{V}$. We input these image-text pairs into the model. The objective is to minimize the cross-entropy loss of predicting the next token. To improve the generalization in various downstream tasks, we adopt a two-stage training pipeline to train our ChartAst-D and ChartAst-S below.
@@
where $\Omega$ is the set of instruction-following data from all tasks in ChartSFT and $\theta$ are the learnable weights initialized from the checkpoint in the Stage I. During training, we sample the data from each task with certain proportions as provided in our experimental setup in Appendix Sec.B. By multitask instructional tuning, our ChartAssistant exhibits strong performance on all the tasks without a task-specific fine-tuning process which is yet a necessary procedure for previous chart-based models [25, 31].
@@ 5.1 Baselines and Evaluation
Baselines. We choose SPHINX [23], ChartLLaMa [7], Unichart [31], Matcha [25], Pix2Struct [19], T5 [36] and Chart-T5 [51] as baselines. ChartLLama and Unichart are the current state-of-the-art models that handles the maximum number of chart tasks and delivers the best overall performance. Besides, Unichart also considers the open-ended QA task. Matcha outperforms previous models in mathematical calculations. Pix2Struct and Donut stands out as an excellent document understanding model. We fine-tune these document models on the train set of the respective evaluation datasets and present the results. T5 is a text-to-text model and needs OCR-based system to extract the data table from the chart image, Chart-T5 is a model modified from T5 for chart-related tasks. We use the results from Unichart [31] for them. Note that ChartAssistant only has undergone task-specific fine-tuning in Chart-to-text [15]. In fact, we can repeat the training set of this data set in the training set to avoid task-specific finetuning, but this will make the generalization ability of the model weaker.
@@
Model aug. human Pew Statista ChartQA OpenCQA MathQA ReferQA
T5 41.0 25.1 10.5 35.3 - 9.3 - -
Chart-T5 74.4 31.8 9.10 37.5 - - - -
Donut 78.1 29.8 7.2 38.2 87.4 13.1 36.3 6.2
SPHINX 11.3 21.7 5.4 16.6 9.4 9.3 4.4 7.2
Pix2Struct 56.0 30.5 10.3 38.0 85.9 12.7 35.6 5.8
Matcha 88.9 38.8 12.2 39.4 89.6 6.5 57.8 8.3
Unichart 87.8 43.9 12.5 38.1 91.1 14.8 23.9 11.9
ChartLLaMa 90.4 48.9 14.2 40.7 90.0 - - -
Ours-D 91.3 45.3 14.1 40.2 92.0 14.9 72.1 64.2
Ours-S 92.0 60.0 15.2 41.0 91.6 15.5 73.9 67.9
@@ 5.2 Main Results
Base type charts. In table 4, we present a comprehensive summary of ChartAssistant’s performance with base type charts across chart-related tasks. It demonstrates that ChartAssistant consistently outperforms the baseline across all tasks with out task-specific fine-tuning. In partivular, we surpass the current leading methods by 9.3% and 1.6% on Human and Augment, respectively. Additionally, in the open-ended question answering, we have 33.1% enhancement relative to Unichart. Besides, we find that most existing models struggle with numerical question answering, while the adaptation of COT answer significantly enhances performance, demonstrating a substantial 16.3% improvement with Matcha. Notably, existing models are currently almost unable to effectively handle the chart referring question answering task. In summation, our model is the top performer across all chart-related tasks without task-specific fine-tuning.
@@
Table 5: A comparison of the results of ChartAssistant with other chart-specific models on five tasks with specialized type charts.
@@
Matcha 17.1 18.3 4.4 7.2 6.3
Unichart 18.4 11.8 5.4 5.9 6.3
Ours-D 68.3 41.1 33.3 42.5 65.2
Ours-S 75.6 42.5 34.5 49.8 68.4
Comparison with foundation model. Despite the strengths of current multimodal models, a domain gap exists between charts and general images, causing difficulty in chart-related tasks. In reaction, we’ve built a small-scale test set for comprehensive evaluation of these models. This test set features five tasks, each with 50 samples containing a variety of chart types. Task samples include both real-world chart-table pairs and tables with API-generated charts. As indicated in table 6, It shows that our 200M model notably outperformed multimodal models like GPT-4V(ision) and Bard, particularly in tasks needing precise chart understanding, such as chart-to-table translation and numerical question answering. Although GPT-4V(ision) does well in chart summarization, we suspect this could be due to the task’s bias towards text fluency, potentially compromising factual accuracy [25].
Table 6: A comparison of the results of ChartAssistant with foundation models on five tasks.
Model C2T Trans. Summ. Open.QA Num.QA Refer.QA
Bard 6.75 20.9 6.2 12.0 22.0
GPT4V 21.6 27.6 9.6 34.0 30.0
Ours-D 80.7 24.9 13.8 66.0 60.0
Ours-S 83.1 26.5 18.2 70.0 66.0

### 15 Feb 2024 v3

@@ Abstract
Charts play a vital role in data visualization, understanding data patterns, and informed decision-making. However, their unique combination of graphical elements (e.g., bars, lines) and textual components (e.g., labels, legends) poses challenges for general-purpose multimodal models. While vision-language models trained on chart data excel in comprehension, they struggle with generalization. To address these challenges, we propose ChartAssistant, a chart-based vision-language model for universal chart comprehension and reasoning. ChartAssistant leverages ChartSFT, a comprehensive dataset covering diverse chart-related tasks with basic (e.g. bars and pies) and specialized (e.g. radars, and bubbles) chart types. It undergoes a two-stage training process, starting with pre-training on chart-to-table parsing to align chart and text, followed by multitask instruction-following fine-tuning. This approach enables ChartAssistant to achieve competitive performance across various chart tasks. Experimental results demonstrate significant performance gains over the state-of-the-art UniChart and Chartllama method, especially outperforming them on real-world chart data with zero-shot setting. The code and data are available at https://github.com/OpenGVLab/ChartAst.
@@ 1 Introduction
Figure 1: A comparison between previous chart-based models and our proposed ChartAssistant. ChartAssistant first aligns the chart and the text by pre-training on the chart-to-table translation task. After performing multitask instruction tuning, it can solve various downstream tasks.
@@
However, chart comprehension is challenging due to the intricate visual marks (e.g. lines, bars and symbols), implicit numerical information, and complex spatial relationships between elements (e.g. axes and labels). Interpreting charts requires specialized knowledge, spatial reasoning, and numerical understanding. The advanced general-purpose multimodal models [48, 21, 47] such as LLaVA [27], trained on natural images, struggle with chart-related tasks due to the specific complexities and relationships unique to charts. Although recent multimodal literate models [29, 19] have achieved impressive results in processing various document-level tasks, they still face difficulties in accurately answering chart-related questions.
@@
In pursuit of universal chart reasoning and comprehension, prior works propose pre-training vision-language models on chart-related tasks as shown in Fig.1(a). For example, both MatCha [25] and UniChart [31] undergo multitask instructional tuning and task-specific fine-tuning, exhibiting good performance on several downstream tasks. However, these models still have severe downsides. Firstly, they fall short in aligning the chart and the associated structured text-form table, which is essential to interpret the relationships between elements in the chart. Although MatCha [25] underscores the importance of chart-text alignment, it presents poor multitask performance due to limited coverage of chart-related tasks. Secondly, the existing training data [32, 30] is deficient in image-text annotations aimed at improving the model’s comprehension of visual elements and mathematical reasoning, as well as annotated data from the specialized chart types such as box-plots. Due to the above factors, existing chart-based models have poor generalization and require task-specific fine-tuning to achieve promising results on various downstream tasks as illustrated in Fig.1(a).
@@
To address these challenges, we propose ChartAssistant, a new multimodal model for universal chart comprehension and reasoning. To improve generalization, ChartAssistant is trained on a large-scale chart-specific instruction-tuning benchmark dubbed ChartSFT. The training process involves a two-stage pre-training pipeline which employs chart-to-table pre-training to align the chart and its structured text and then perform joint tuning on multiple chart-related tasks as shown in Fig.1(b). As a result, our ChartAssistant can achieve good results on various chart-related tasks with a single model. We implement ChartAssistant with two variants, i.e. ChartAst-D and ChartAst-S. ChartAst-D is built upon Donut [15], a lightweight ($260$M parameters) but powerful vision-language model for visual document understanding. While ChartAst-S is built upon SPHINX [23], a large ($13$B parameters) vision-language model for universal multimodal comprehension. Inherited from SPHINX, our ChartAst-S obtains enhanced chart representation by dynamic resolution processing and mixed visual encoders. Therefore, ChartAst-S offers increased robustness and usability for chart understanding, demonstrating strong performance in various chart-related tasks.
@@
Before conducting multitask instruction tuning, as done in existing research [31, 25], we start with pre-training ChartAssistant on the chart-to-table translation task as shown in Fig.1(b). This task involves parsing a chart and generating a Markdown table. It shares similarities with dense captioning for natural images, allowing the model to interpret the elements and relationships within the chart. Similar to the role of image captioning in training multimodal models [27, 38, 45], chart-to-table translation facilitates alignment between the chart and its structured text. Following pre-training, we proceed with multitask instruction tuning using ChartSFT. This two-stage training approach enables ChartAssistant (a single model) to achieve strong performance across a range of chart-related tasks.
The contributions of this paper can be summarized as follows. 1) We present ChartAssistant, a vision-language model for chart comprehension and reasoning. ChartAssistant is versatile enough to solve various chart-related tasks across a wide range of chart types. 2) We build a chart-specific visual instruction-following benchmark dubbed ChartSFT. ChartSFT surpasses existing chart-based benchmarks with its larger instruction-following data corpus, a broader range of tasks and chart types, and more comprehensive data annotations. 3) Extensive experimental results on various downstream tasks demonstrate that ChartAssistant surpasses the previous SoTA method UniChart [31] by 50.0%, 28.1% performance gain on numerical QA and ChartQA, respectively. Notably, ChartAssistant continues to significantly outperform existing Chart-related models in the zero-shot setting, with 29.5% performance gain on RealCQA [2] compared with Unichart and 23.6% performance gain on ChartLLM [18] compared with Chartllama [6].
@@ 2.2 Chart-specific Vision-Language Model
Some methods modify vision-language models for chart-related tasks [6, 26] or develop plugin for LLM to understand the chart [44]. Matcha [25] extends Pix2Struct [19] by integrating mathematical reasoning and chart data extraction tasks, excelling at chart question answering and chart summarization. Unichart [31] undergoes multitask instruction tuning for more chart-related tasks, establishing itself as the most versatile and effective chart vision-language model currently available. However, these methods have limitations. Furthermore, these models struggle with mathematical computations, limiting their effectiveness and range of applicable chart types.
Contrastingly, we propose ChartSFT, the most extensive dataset to date, supporting a wide variety of chart tasks and types. We develop ChartAssistant using ChartSFT with a two-stage training strategy, capable of handling diverse chart-related tasks.
@@ 3 ChartSFT
We construct a large-scale chart-specific instruction-tuning benchmark called ChartSFT by collecting data from various tasks. The composition of ChartSFT is shown in Table 1, as extensively described below. Our ChartSFT consists of $39$M pieces of chart-text annotated data, $4.75$ and $5.62$ times larger than MatCha [25] and UniChart [31], respectively, as illustrated in Fig.3. ChartSFT contains charts with both base and specialized types, as presented in Sec. 3.1 and Sec. 3.2, respectively.
Overall, our ChartSFT encompasses nine types of charts by collecting data from various sources as shown in table 10. First, most charts with base types including bar, line, dot-line, and pie are collected from several existing datasets [30, 32, 13, 37, 22, 39, 14]. Second, we also generate some charts with base types from arXiv tables [1] and data augmentation techniques (e.g. various APIs and figure parameters). In particular, we use ChatGPT to suggest the proper chart type given each table data from arXiv. Third, we synthesize table data which is appropriate for depicting charts with specialized types.
@@ 3.2 Charts with Specialized Types
Previous chart-based models have exhibited poor performance when dealing with specialized chart types, such as radar, area, histogram, bubble, and box-plot. To enhance the model’s generalization capabilities, we have trained our ChartAssistant on these charts with specialized types. To overcome the challenge of obtaining large-scale real-world chart data, we have employed synthetic data generation techniques. For more detailed information, please refer to Appendix Sec. A.1. Through this approach, we can obtain a substantial and diverse collection of complex charts across these specialized types.
@@ 4.1 Architecture
The key to completing the tasks related to charts lies in accurately understanding the content of the charts. As shown in Fig. 4, we implement ChartAssistant with two variants, i.e. ChartAst-D and ChartAst-S, which have 260M and 13B parameters in total. In addition, their input image resolutions are $224\times 224$ and $448\times 448$, respectively. Both ChartAst-D and ChartAst-S perform well in many chart-related tasks. But ChartAst-D has a smaller size and ChartAst-S enjoys better generalization.
@@
ChartAst-S is a large vision-language model for chart understanding built upon Sphinx [23]. For high-resolution images, it preserves the original information through sampling and partitioning methods, ensuring greater fidelity to the image content. Moreover, Sphinx leverages the abundant prior knowledge of LLM [40] to handle various tasks such as visual question answering and image summarization. Specifically, for an input image $X_{V}$. ChartAst-S incorporates multiple visual encoders to extract more informative visual features $Z_{V}$, such as DINOv2 [33], CLIP [35], and ConvNeXt [43]. Unlike ChartAst-D where visual tokens are involved in a language decoder with a cross-attention module, ChartAst-S directly appends visual tokens to the text tokens $X_{q}$. The merged tokens are then fed into the LLM to generate the response. Thanks to the intricate design of the visual encoder and the powerful reasoning ability of LLM, ChartAst-D generalizes well in various real-world chart-related applications.
@@ 4.2 Training
In our ChartSFT, we have a corresponding instruction $X_{q}$ and response $Y_{q}$ for each image $X_{V}$. We input these image-text pairs into the model. The objective is to minimize the cross-entropy loss of predicting the next token. To improve the generalization in various downstream tasks, we adopt a two-stage training pipeline to train our ChartAst-D and ChartAst-S below. Charts are special images that visualize the data and underlying relationships between elements in the chart. Understanding the numerical values and their meanings is a prerequisite for completing downstream tasks related to charts. Therefore, we employ Chart-to-Table translation as a pre-training task, aiming to enable the model to understand the correspondence between charts and tables, which has also been utilized as a part of a pre-training task in MMC [26] and Matcha [25].
@@
where $\Omega$ is the set of instruction-following data from all tasks in ChartSFT and $\theta$ are the learnable weights initialized from the checkpoint in the Stage I. During training, we sample the data from each task with certain proportions as provided in our experimental setup in Appendix Sec.B. By multitask instructional tuning, our ChartAssistant exhibits strong performance on all the tasks.
@@ 5.1 Baselines and Evaluation
Baselines. We choose SPHINX [23], Blip2-flant5-xl [21], Qwen-VL [3], ChartLLaMa [6], Unichart [31], Matcha [25], Pix2Struct [19], T5 [36] and Chart-T5 [50] as baselines. ChartLLama and Unichart are the current state-of-the-art models that handles the maximum number of chart tasks and delivers the best overall performance. Besides, Unichart also considers the open-ended QA task. Matcha outperforms previous models in mathematical calculations. Pix2Struct and Donut stands out as an excellent document understanding model. We fine-tune these document models on the train set of the respective evaluation datasets and present the results. T5 is a text-to-text model and needs OCR-based system to extract the data table from the chart image, Chart-T5 is a model modified from T5 for chart-related tasks. We use the results from Unichart [31] for them. SPHINX [23], Blip2[21] and Qwen-VL [3] are all commonly used large vision-language models at present. We observe that these models underperform in processing Chart tasks. Finally, Chartllama, utilizing LLaVA for training on Chart data, demonstrates superior performance in Chart tasks. Therefore, we only compare with Chartllama.
@@
Model Size aug. human Pew Statista ChartQA OpenCQA MathQA ReferQA
T5 223M 41.0 25.1 10.5 35.3 - 9.3 - -
Chart-T5 400M 74.4 31.8 9.10 37.5 - - - -
Donut 260M 78.1 29.8 7.2 38.2 87.4 13.1 36.3 6.2
Pix2Struct 300M 81.6 30.5 10.3 38.0 85.9 12.7 35.6 5.8
SPHINX 13B 11.3 21.7 3.2 4.1 9.4 5.9 4.4 7.2
Qwen 9.6B 78.9 44.3 0.5 2.6 - 1.3 4.8 4.9
Blip2 4B 1.4 7.8 0.2 0.8 - 1.7 6.4 0.4
Matcha 300M 88.9 38.8 12.2 39.4 89.6 6.5 57.8 8.3
Unichart 260M 87.8 43.9 12.5 38.1 91.1 14.8 23.9 11.9
ChartLLaMa 13B 90.4 48.9 14.2 40.7 90.0 4.7 5.8 9.9
ChartAst-D 260M 91.3 45.3 14.0 40.2 92.0 14.9 72.1 64.2
ChartAst-S 13B 93.9 65.9 15.2 41.0 91.6 15.5 73.9 67.9
@@ 5.2 Main Results
Base type charts. In table 4, we present a comprehensive summary of ChartAssistant’s performance with base type charts across chart-related tasks. It demonstrates that ChartAssistant consistently outperforms the baseline across all tasks. In partivular, we surpass the current leading methods by 17% and 2.5% on ChartQA-human and ChartQA-augment, respectively. Besides, we find that most existing models struggle with numerical question answering, while the adaptation of COT answer significantly enhances performance, demonstrating a substantial 16.1% improvement with Matcha. Notably, existing models are currently almost unable to effectively handle the chart referring question answering task. At last, In summation, our model is the top performer across all chart-related tasks. It is important to note that both Unichart and Matcha’s results are given after fine-tuning on the training set of the test dataset, whereas ChartAssistant’s results are obtained using a single model after training is complete, except for Chart-to-Text, because this dataset contain too few groundtruth references, which means that the results must be very close to the reference targets to achieve high scores when evaluating using BLEU-4 [6].
@@
Table 5: A comparison of the results of ChartAssistant with other chart-specific models on five tasks with specialized type charts. Use BLEU to evaluate summarization and open-ended QAs.
@@
Matcha 17.1 6.3 5.1 7.2 -
Unichart 18.4 6.3 5.4 5.9 -
Chartllama 19.4 9.2 8.4 2.4 -
ChartAst-D 68.3 19.7 25.7 42.5 65.2
ChartAst-S 75.6 22.0 27.8 49.8 68.4
6 Zero-shot Study
In addition to outperforming the current best methods on common datasets such as ChartQA and Chart-to-text, ChartAssistant-S demonstrates its excellence. To validate the model’s generalizability, it is necessary to test on samples not included in the training set. For this purpose, we have collected data from StructChart [44], RealCQA [2], and ChartLLM [18] for tasks like chart-to-table translation, chart-based question answering, and summarization. The results indicate that ChartAssistant exhibits superior zero-shot performance across all tasks, surpassing current methods. We also present a selection of comparative examples in the supplementary materials for visualization. For evaluation, RealCQA uses accuracy within a 5% error margin, ChartLLM employs GPT-4 scoring used in Chartllama [6], while StructChart is evaluated using $RMS_{F_{1}}$ metrics. As shown in table 6, we find that in the zero-shot setting, Chartllama performs poorly in precise numerical question answering but excels in summarization tasks. We attribute this to the robust language capabilities of LLM. On the other hand, ChartAssistant surpasses existing models in tasks such as precise numerical question answering in OCR and summarization, which involves generating long texts. Furthermore, we observe that if the model’s decoder is not powerful enough, errors are more likely to occur in the zero-shot setting when tasked with generating long text outputs, such as in summarization or providing answers in COT format. The use of Large Language Models (LLMs) can significantly alleviate this issue.
Table 6: In comparison with other chart-related multimodal models in a zero-shot setting, ChartAssistant-S significantly outperforms existing models across all tasks in the zero-shot scenario.
RealQA
Model Math Extract ChartLLM StructChart
Unichart 13.0 33.0 11 41.5
Matcha 16.0 27.5 11 23.3
Chartllama 10.0 13.0 55 38.3
ChartAst-D 15.0 36.0 13 39.4
ChartAst-S 32.0 43.5 68 45.3